OpenStudy (anonymous):

write the wquation of the line which contains the point (-2,5) and has a slope of 4, in slope intercept form

OpenStudy (bmorse):

y-5=4(x-(-2)) solve for y y=5+4x+8 y=4x+13

OpenStudy (anonymous):

that's not in slope intercept form

OpenStudy (anonymous):

what about write the equation y=1/4x-4 in standard form

OpenStudy (bmorse):

Multiply both sides by 4 to get rid of the fraction 4(y=1/4x-4) 4y=x-16 x-4y=16